# Engineering Software 2.0 by Interpolating Neural Networks: Unifying Training, Solving, and Calibration

### v1

@@ 1 Introduction
The evolution of software programming methodologies has transitioned from reliance on rigidly hard-coded instructions, governed by explicitly defined human rules, to the adoption of neural network-based algorithms. The transition is coined as "from Software 1.0 to Software 2.0" [1]. The shift towards Software 2.0 has facilitated significant advancements in the domain of large language models (LLMs) and other foundational models [2, 3]. However, the application of these technologies within the field of engineering and science introduces distinct challenges. These challenges include scarcity and multi-modal aspect of data, low generalization error compared to classical numerical methods [4, 5, 6], slow inference (function evaluation) time that challenges real-time control [7, 8, 9, 10] and statistical modeling, and resource-heavy model architecture [11, 12, 13, 14].
@@ 2 INN: Unifying Training/Learning, Solving, and Calibration
Interpolating Neural Network (INN) is a systematically pruned neural network based on interpolation theories [17, 22] that leverage classical numerical methods such as finite element [23] and meshfree methods [24, 25]. Although these numerical methods and the neural network theory have been developed by seemingly different groups of researchers, they share one thing in common; the goal is to find a function. The former finds a solution field of a differential equation while the latter finds an input-output relationship. INN is built upon this similarity and it forms the backbone of Engineering Software 2.0 that unifies training, solving, and calibration, as illustrated in Figure 1a.
@@ 3 Benchmark Demonstration
To demonstrate the advantages of INN for training, calibration, and solving, we present several benchmarks for each subject. The benchmark problems include training of data that follow scientific equations, classification of spiral data, classification of the fashion-MNIST dataset [33], calibration of parameters in the scientific equations and electrode manufacturing data [38], and solving of a parameterized partial differential equation (PDE). Except for the fashion-MNIST classification, all benchmark demonstrations are provided in the supplementary information, clearly revealing significant benefits such as less trainable parameters, high accuracy, less memory requirement, fast training, and inference time. One can also find how the INN hyperparameters are selected in Supplementary Information.

### v2

@@
Engineering Software 2.0 by Interpolating Neural Networks: Unifying Training, Solving, and Calibration
Chanwook Park
Department of Mechanical Engineering
Northwestern University
Evanston, IL 60208
chanwookpark2024@u.northwestern.edu
&Sourav Saha
Theoretical and Applied Mechanics Program
Northwestern University
Evanston, IL 60208
sourav023@u.northwestern.edu
&Jiachen Guo
Theoretical and Applied Mechanics Program
Northwestern University
Evanston, IL 60208
jiachenguo2020@u.northwestern.edu
&Xiaoyu Xie
Department of Mechanical Engineering
Northwestern University
Evanston, IL 60208
xiaoyuxie2020@u.northwestern.edu
&Satyajit Mojumder
Theoretical and Applied Mechanics Program
Northwestern University
Evanston, IL 60208
satyajitmojumder2022@u.northwestern.edu
&Yangfan Li
Department of Mechanical Engineering
Northwestern University
Evanston, IL 60208
YangfanLi2024@u.northwestern.edu
&Stefan Knapik
Department of Mechanical Engineering
Northwestern University
Evanston, IL 60208
stefan.knapik@u.northwestern.edu
&Miguel A. Bessa
School of Engineering
Brown University
Providence, RI
miguel_bessa@brown.edu
&Dong Qian
Department of Mechanical Engineering
University of Texas at Dallas
Richardson, TX 75080
dong.qian@utdallas.edu
&Wei Chen
Department of Mechanical Engineering
Northwestern University
Evanston, IL 60208
weichen@northwestern.edu
&Gregory J. Wagner
Department of Mechanical Engineering
Northwestern University
Evanston, IL 60208
gregory.wagner@northwestern.edu
&Jian Cao
Department of Mechanical Engineering
Northwestern University
Evanston, IL 60208
jcao@northwestern.edu
&Wing Kam Liu
Department of Mechanical Engineering
Northwestern University
Evanston, IL 60208
w-liu@northwestern.edu
Equal contribution with Coauthor 1 and Coauthor 2Corresponding author
@@ 1 Introduction
The evolution of software programming methodologies has transitioned from reliance on rigidly hard-coded instructions, governed by explicitly defined human rules, to the adoption of neural network-based algorithms. The transition is coined as "from Software 1.0 to Software 2.0" [1]. The shift towards Software 2.0 has facilitated significant advancements in the domain of large language models (LLMs) and other foundational models [2, 3]. However, the application of these technologies within the field of engineering and science introduces distinct challenges. These challenges include scarcity and multi-modal aspect of data, low generalization error compared to classical numerical methods [4, 5, 6, 7], slow inference (function evaluation) time that challenges real-time control [8, 9, 10, 11] and statistical modeling, and resource-heavy model architecture [12, 13, 14, 15].
@@ 2 INN: Unifying Training/Learning, Solving, and Calibration
Interpolating Neural Network (INN) is a systematically pruned neural network based on interpolation theories [18, 7] that leverage classical numerical methods such as finite element [23, 24, 25] and meshfree methods [26, 27]. Although these numerical methods and the neural network theory have been developed by seemingly different groups of researchers, they share one thing in common; the goal is to find a function. The former finds a solution field of a differential equation while the latter finds an input-output relationship. INN is built upon this similarity and it forms the backbone of Engineering Software 2.0 that unifies training, solving, and calibration, as illustrated in Figure 1a.
@@ 3 Benchmark Demonstration
To demonstrate the advantages of INN for training, calibration, and solving, we present several benchmarks for each subject. The benchmark problems include training of data that follow scientific equations, classification of spiral data, classification of the fashion-MNIST dataset [35], calibration of parameters in the scientific equations and electrode manufacturing data [40], and solving of a parameterized partial differential equation (PDE). Except for the fashion-MNIST classification, all benchmark demonstrations are provided in the supplementary information, clearly revealing benefits of INN such as less trainable parameters, high accuracy, less memory requirement, and fast training and inference time. One can also find how the INN hyperparameters are selected.